A caterer charges a fixed cost for preparing dinner plus an additional cost for each person served. You know the cost for 100 students will be $750 an the cost for 150 students will be $1050
Find the caterers fixed cost And the cost per student served.

Answer by checkley77(12844)

X+100Y=750
X+150Y=1,050 SUBTRACT.
----------------
-50Y=-300
Y=-300/-50
Y=$6 PER STUDENT.
X+100*6=750
X+600=750
X=750-600
X=$150 FIXED COST.
PROOF:
PROOF:
150+150*6=1,050
150+900=1,050
1,050=1,050